OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cosine of five pi divided by twelve.

OpenStudy (jdoe0001):

\(\bf cos\left(\frac{5\pi}{12}\right)\\ \textit{keep in mind that } \frac{1}{12}\times 2 = \frac{1}{6}\\ cos\left(\frac{5\pi}{12}\right)\implies cos\left(\cfrac{ \frac{5\pi}{6} }{ 2 }\right)\) so use the half-angle identity for the cosine then